OpenStudy (anonymous):

solve: \[e ^{x-4}=9\] iv got x=5.197 but i think its wrong

OpenStudy (zzr0ck3r):

ln of both sides ln(e^(x-4)) = ln(9) (x-4)*ln(e) = ln(9) x-4 = ln(9) x = ln(9)+4

OpenStudy (anonymous):

oh okay, thanks! @zzr0ck3r